An investor has $8000 invested at 9% and wishes to invest some addtional money at 16% in order to return at least 12% on total investment. What is the least amount of money that should be invested?
Answer by alicealc(293)

let the additional money = m
8000*9/100 + m*16/100 = (8000+m)*12/100
72000/100 + 16m/100 = 96000/100 + 12m/100
720 + 16m/100 = 960 + 12m/100
16m/100 - 12m/100 = 960 - 720
4m/100 = 240
4m = 240*100
4m = 24000
m = 24000/4
m = 6000
so,the least amount of additional money should be invested is $6000